# High dimensional hyperbolic Coxeter groups that virtually fiber

### v1

@@ 5.1 Topology of links, spheres, and balls
Finally, we check that the actual simplicial complexes coincide. For the simplicial complex $\operatorname{lk}(\square,X)$, the set of vertices $\{\square^{\prime}_{0},\ldots,\square^{\prime}_{k}\}$ span a $k$-simplex if and only if the cubes $\{\square^{\prime}_{0},\ldots,\square^{\prime}_{k}\}$ are all contained in common larger cube $\square^{\prime\prime}$. Considering $\operatorname{lk}(v,X)$, this collection of cubes correspond to a collection of simplices $\{\tau^{\prime}_{0},\ldots,\tau^{\prime}_{k}\}$ in $\operatorname{lk}(v,X)$ all containing $\tau$ as a codimension one face. The larger cube $\square^{\prime\prime}$ leads to a larger simplex $\tau^{\prime\prime}\subset\operatorname{lk}(v,X)$ containing all of the $\tau^{\prime}_{i}$ as facets. Viewing the $\tau^{\prime}_{i}$ as a collection of vertices in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$, it follows that they span a simplex in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$. This shows that simplices in $\operatorname{lk}(\square,X)$ give rise to simplices in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$. One can reverse this argument to show the converse. Thus the bijection on vertex sets described above also bijectively identifies simplices of the two complexes, establishing that the two simplicial complexes are isomorphic. ∎
@@
Next we turn our attention to the topology of the spherical sets $S(\sigma,\Sigma)$. Recall that this set is a small $\epsilon$-sphere centered around the simplex $\sigma$, and it inherits a natural polyhedral structure from the ambient simplicial complex $\Sigma$. Indeed, each point $p\in S(\sigma,\Sigma)$ is contained in the interior of a unique simplex in $\Sigma$ that splits as a join $\sigma_{p}*\tau_{p}$ where $\sigma_{p}\subset\sigma$ is a face of $\sigma$ and $\tau_{p}$ is a simplex disjoint from $\sigma$. The intersection of $S(\sigma,\Sigma)$ with the simplex $\sigma_{p}*\tau_{p}$ is a copy of the polyhedron $\sigma_{p}\times\tau_{p}$, which we can think of as the “$\epsilon$-slice” $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$ in the join:
@@
Now given $\tau$, express it as a join $v*\tau^{\prime}$ where $v\in\hat{S}(\sigma,\Sigma)$ is a simplex and $\tau^{\prime}\subset\hat{S}(\sigma,\Sigma)$ is a codimension one face. By the inductive hypothesis, there exists a vertex $w^{\prime}\subset\sigma$ with the property that $w^{\prime}*\tau^{\prime}$ is a simplex in $\Sigma$. If $v$ is connected to $w^{\prime}$, then $\tau*w^{\prime}$ is a simplex since $\Sigma$ is flag, and we are done. If $v$ is not connected to $w^{\prime}$, then $v$ is connected to some vertex $v^{\prime}\in\sigma$, with $v^{\prime}\neq w^{\prime}$. For each vertex $w\in\tau^{\prime}$, consider the ordered $4$-tuple $\{v,v^{\prime},w^{\prime},w\}$. This defines a square in $\Sigma$, and we know by construction that $v$ is not connected to $w^{\prime}$. Then $5$-largeness forces $w$ to be connected to $v^{\prime}$. Since this holds for arbitrary $w\in\tau^{\prime}$, we conclude that all the vertices of $\tau^{\prime}$ are connected to $v^{\prime}$. Again, the fact that $\Sigma$ is flag implies that $\tau*v^{\prime}$ is a simplex, thus completing the proof. ∎
@@
$\sigma$$\hat{B}(\sigma,\Sigma)$$\sigma$$S(\sigma,\Sigma)$$\sigma$$\hat{S}(\sigma,\Sigma)$
@@
We first consider the topology of $\hat{S}(\sigma,\Sigma)$. There is a natural map $\rho:S(\sigma,\Sigma)\rightarrow\hat{S}(\sigma,\Sigma)$, given by projecting away from the simplex $\sigma$. More explicitly, on each simplex
@@
incident to $\sigma$, the intersection with $S(\sigma,\Sigma)$ is the polyhedral slice $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$, while the intersection with $\hat{S}(\sigma,\Sigma)$ is the top slice $\tau_{p}\cong(\sigma_{p}\times\tau_{p}\times\{1\})/\sim$. The map $\rho$ is then just the projection to $\{1\}$ in the interval coordinate. In terms of the polyhedral structure on $\hat{S}(\sigma,\Sigma)$, $\rho$ collapses the $\sigma_{p}$ factor in each polyhedron $\sigma_{p}\times\tau_{p}$. Note that $\rho$ is surjective by 5.3. Thus the subset of the preimage under $\rho$ of any point in any given polyhedron of $S(\sigma,\Sigma)$ is contractible or empty.
@@
To verify the corresponding statement for $\hat{B}(\sigma,\Sigma)$, we note that $\hat{B}(\sigma,\Sigma)$ is the union of simplices of the form $\sigma_{p}*\tau_{p}$, where $\tau_{p}$ is a face in $\hat{S}(\sigma,\Sigma)$ and $\sigma_{p}$ is a face of $\sigma$ by 5.3. We now form an auxiliary space $B(\sigma,\Sigma)$ and a continuous quotient map $\phi:B(\sigma,\Sigma)\rightarrow\hat{B}(\sigma,\Sigma)$. Recall that the space $S(\sigma,\Sigma)$ is the $\epsilon$-sphere centered at $\sigma$, identified with the $\epsilon$-slice $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$ in the join description of each simplex incident to $\sigma$. We let $B(\sigma,\Sigma)$ denote the closed $\epsilon$-ball centered at $\sigma$. Thus $B(\sigma,\Sigma)$ gets identified with $(\sigma_{p}\times\tau_{p}\times[0,\epsilon])/\sim$ in the join description of each simplex incident to $\sigma$; recall that at the 00-level we collapse the $\tau_{p}$ coordinates to points. Since we can contract the interval factors down, we clearly have that $B(\sigma,\Sigma)$ deformation retracts to $\sigma$, hence it is contractible.
The space $B(\sigma,\Sigma)$ contains a copy of $S(\sigma,\Sigma)$ at level $\epsilon$, and we now define $\phi$ to be the quotient map $\rho$ on the subset $S(\sigma,\Sigma)$. As we can stretch out the $[0,\epsilon]$ coordinate to $[0,1]$, we see that the quotient can naturally be identified with $\hat{B}(\sigma,\Sigma)$. Moreover point preimages under $\phi$ are either points when $x\notin\hat{S}(\sigma,\Sigma)$ or coincide with the point preimage under $\rho$ when $x\in\hat{S}(\sigma,\Sigma)$), and hence point preimages are always contractible. Thus the map $\phi$ is cell-like, and hence it is a homotopy equivalence. We conclude that the space $\hat{B}(\sigma,\Sigma)$ is contractible, as desired. ∎
@@ 5.2 Analysis of combinatorial spheres in thickenings
We now return to analyzing the thickenings $T_{n}$ from Section 4 in order to prove 4.1. Adopting all notation from that section, recall that the goal is to prove that $\operatorname{cd}(S(\sigma,T_{n}))\leq n-1$ for all nonempty simplices $\sigma$ of $T_{n}$. Equivalently, we want to show that $H^{i}(S(\sigma,T_{n}))=0$ for all $i\geq n$ and $\emptyset\neq\sigma\subset T_{n}$.
@@
To complete the proof of the lemma, it remains to verify property (3). Let $Z$ denote the union of all the cubes containing $\square_{\sigma}$ and $\hat{Z}$ be the cube subcomplex of $Z$ consisting of all cubes in $Z$ that do not intersect $\square_{\sigma}$. There is a natural homeomorphism $\hat{Z}\cong\square_{\sigma}\times\operatorname{lk}(\square_{\sigma},X)$, so $\hat{Z}$ is homotopy equivalent to $\operatorname{lk}(\square_{\sigma},X)$. From the definition of $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$, any simplex $\tau\subset\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ has the property that $\alpha(V(\tau))$ is disjoint from $\square_{\sigma}$, and hence it lies in $V(\hat{Z})$. In other words, the entire complex $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ lies “above” the cube complex $\hat{Z}$.
Now consider the thickening $\operatorname{Th}_{\alpha}(\hat{Z})$ obtained by the restriction $\alpha:\alpha^{-1}(V(\hat{Z}))\rightarrow V(\hat{Z})$. This is a subcomplex of $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ that is homotopy equivalent to $\hat{Z}$, and hence to $\operatorname{lk}(\square_{\sigma},X)$. To complete the verification of property (3) it therefore suffices to show that $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ deformation retracts to $\operatorname{Th}_{\alpha}(\hat{Z})$.
@@
Recall that $\hat{Z}$ is defined as a subcomplex of the cube complex $Z$. To construct the deformation retraction, we proceed by induction on the dimension of the cubes in $Z$. More precisely, we will consider a cube $\square\subset Z$ containing $\square_{\sigma}$ and the restriction of the two complexes $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ and $\operatorname{Th}_{\alpha}(\hat{Z})$ lying above $\hat{Z}\cap\square$, i.e., whose vertices $\alpha$-project to vertices in $\hat{Z}\cap\square$. We induct on $\dim(\square)$ and show that in each dimension we can choose a deformation retraction compatible with the previously-constructed deformation retractions. This will allow the deformation retractions associated with each cube to glue compatibly and yield a global deformation retraction of $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ to $\operatorname{Th}_{\alpha}(\hat{Z})$.
@@
The bottom of the induction is the case of a cube $\square\subset Z$ with $\dim(\square)=\dim(\square_{\sigma})+1$. In this case, $\square=\square_{\sigma}\times I$, and $\hat{Z}\cap\square$ is the face of $\square$ opposite $\square_{\sigma}$. This forces the two subcomplexes above $\hat{Z}\cap\square$ to coincide, and hence we can take the identity map as the deformation retraction. Now for the inductive step, we assume that we are already provided the deformation retractions for the portions of the simplex that lies above the relevant lower dimensional faces of the cube $\hat{Z}\cap\square$. We would like to extend the deformation retraction to the rest of the simplicial complex lying above $\hat{Z}\cap\square$.
@@
On the other hand, the subcomplex $\operatorname{Th}_{\alpha}(\hat{Z}\cap\square)$ is the $\alpha$-thickening of the cube complex $\hat{Z}\cap\square$. However this cube complex is spanned by the vertices $V(\square)\setminus V(\square_{\sigma})$, and thus is homotopic to the contractible set $\partial\square\setminus\square_{\sigma}$. We conclude that $\operatorname{Th}_{\alpha}(\hat{Z}\cap\square)$ is also a contractible subset of $\Delta$. Since both spaces are contractible, the relative homotopy groups all vanish and the desired deformation retraction exists. This completes the inductive step and verifies statement (3), completing the proof of the lemma. ∎
@@
$\operatorname{cd}(\operatorname{lk}(\sigma,T_{n}))=\operatorname{cd}(\operatorname{lk}(\tau,T_{n-1}))\leq n-2.$
@@
Note that every vertex $v$ in $\hat{S}(\sigma,T_{n})$ is at distance one from some vertex of $\sigma$ by definition. For such a vertex $v$, we can look at all the vertices of $\sigma$ that are adjacent to $v$. This defines a face of $\sigma$, and gives us a map $\pi:V(\hat{S}(\sigma,T_{n}))\rightarrow\mathcal{F}(\sigma)$ from the set of vertices in $\hat{S}(\sigma,T_{n})$ to the set $\mathcal{F}(\sigma)$ of faces of $\sigma$. We think of $\pi(v)$ as encoding the face of $\sigma$ that is closest to the vertex $v$. We can now partition the vertices of $\hat{S}(\sigma,T_{n})$ according to their image under $\pi$; see Figure 8. Next, use this partition to form a filtration of $\hat{S}(\sigma,T_{n})$ by subcomplexes. Denote by $S_{i}$ the subcomplex of $\hat{S}(\sigma,T_{n})$ spanned by the preimage under $\pi$ of all faces of codimension $\leq i$. We clearly have containments $S_{0}\subset S_{1}\subset\cdots\subset S_{k}=\hat{S}(\sigma,T_{n})$, where $k=\dim(\sigma)$. This gives us a filtration of the complex $\hat{S}(\sigma,T_{n})$ by the subcomplexes $S_{i}$; again, see Figure 8.
$\sigma$codim-00 faces of $\sigma$$\sigma$$\pi$$\pi$$\pi$codim-$1$ faces of $\sigma$$\sigma$$\pi$$\pi$$\pi$codim-$2$ faces of $\sigma$
$\sigma$$S_{0}$$\sigma$$S_{1}$$\sigma$$S_{2}=\hat{S}(\sigma,T_{n})$
Figure 8: The top row shows $\pi:V(\hat{S}(\sigma,T_{n}))\rightarrow\mathcal{F}(\sigma)$ for faces of $\sigma$ in each codimension. The bottom row shows the filtration $S_{0}\subset S_{1}\subset\cdots\subset\hat{S}(\sigma,T_{n})$.
@@
By construction, the vertices in $L_{\tau}$ can be identified with vertices in $\operatorname{lk}(\tau,T_{n})$ by viewing the link as a subcomplex in the flag complex $T_{n}$. We claim that the vertices of ${\operatorname{lk}(\tau,T_{n})\setminus\{L_{\tau}\cup\hat{\sigma}\}}$ are precisely the vertices of $\operatorname{lk}(\tau,T_{n})$ at combinatorial distance one from $\hat{\sigma}$. Indeed, if $v$ is at combinatorial distance one from $\hat{\sigma}$, this means there is some vertex $w\in\hat{\sigma}$ connected by an edge to $v$. In terms of the embedding of $\operatorname{lk}(\tau,T_{n})$ into $T_{n}$, the simplex $\hat{\sigma}\subset\operatorname{lk}(\tau,T_{n})$ corresponds to the face of $\sigma$ opposite $\tau$, so $w\in\sigma\setminus\tau$. Since $v\in\operatorname{lk}(\tau,T_{n})$ is connected by an edge to $w$, we have that $\tau\cup\{w\}\subseteq\pi(v)$. This forces the codimension of $\pi(v)$ to be at most $i-1$, which means that $v\in S_{i-1}$. This argument can be reversed to show that the vertices in $L_{\tau}$ are precisely the vertices in $\operatorname{lk}(\tau,T_{n})$ whose combinatorial distance to $\hat{\sigma}$ is at least two. Of course, the vertices of $\hat{\sigma}$ in $\operatorname{lk}(\tau,T_{n})$ are precisely those in the face of $\sigma$ opposite $\tau$, so they do not lie in $S_{i-1}$. We conclude that the set $\partial L_{\tau}$ consists of precisely the vertices in $\operatorname{lk}(\tau,T_{n})$ that are exactly at distance $1$ from $\hat{\sigma}$.
@@
Consider a pair of simplicial complexes $Z,Z^{\prime}$ that can be expressed as gluings $Z=A\cup_{C}B$ and $Z^{\prime}=A\cup_{C}B^{\prime}$, where each inclusion of $C$ into $A$ induces the same maps on cohomology. Assume $B$ is contractible, $\operatorname{cd}(Z)\leq n-1$, and $\operatorname{cd}(B^{\prime})\leq n-1$. Then $\operatorname{cd}(Z^{\prime})\leq n-1$.
@@
$\cdots\longrightarrow H^{i}(Z)\longrightarrow H^{i}(A)\longrightarrow H^{i}(C)\longrightarrow H^{i+1}(Z)\longrightarrow\cdots$
We also know by hypothesis that $\operatorname{cd}(Z)\leq n-1$, so $H^{i}(Z)=0$ for $i\geq n$. The long exact sequence now tells us that the map $H^{i}(A)\rightarrow H^{i}(C)$ induced by the inclusion $C\hookrightarrow A$ is an isomorphism when $i\geq n$ and a surjection when $i=n-1$.
@@
$\cdots\longrightarrow H^{i-1}(C)\longrightarrow H^{i}(Z^{\prime})\longrightarrow H^{i}(A)\oplus H^{i}(B^{\prime})\longrightarrow H^{i}(C)\longrightarrow\cdots$
By hypothesis, we have that $\operatorname{cd}(B^{\prime})\leq n-1$, so $H^{i}(B^{\prime})=0$ for $i\geq n$. In particular, when $i\geq n$ the sequence above reduces to:
$\cdots\longrightarrow H^{i-1}(C)\longrightarrow H^{i}(Z^{\prime})\longrightarrow H^{i}(A)\longrightarrow H^{i}(C)\longrightarrow\cdots$
From our earlier analysis, we also know that the map $H^{i}(A)\rightarrow H^{i}(C)$ is an isomorphism when $i\geq n$, and a surjection when $i=n-1$. This forces $H^{i}(Z^{\prime})=0$ in the range $i\geq n$, which proves the lemma. ∎
@@
We want to show that $\operatorname{cd}(S(\sigma,T_{n}))\leq n-1$ for every nonempty simplex $\sigma\subset T_{n}$, . From 5.4, we have a homotopy equivalence $S(\sigma,T_{n})\simeq\hat{S}(\sigma,T_{n})$, so we instead show that $\operatorname{cd}(\hat{S}(\sigma,T_{n}))\leq n-1$. As discussed above, the set $\hat{S}(\sigma,T_{n})$ has a filtration
@@
We will consider the cohomological dimension of each stage in this filtration, and check that it always remains less than or equal to $n-1$.
@@
  • $B$ is the subcomplex of $Z$ consisting of all simplices incident to $\hat{\sigma}$ (from 5.10), which coincides with $\hat{B}(\hat{\sigma},Z)$, and is thus contractible by 5.4, since $T_{n}$, hence $Z$, is flag and $5$-large;
@@
  • $C\subseteq S_{i-1}$ from the conclusions about $\partial L_{\tau}$ in 5.10, hence $C$ is contained in $B$ and $B^{\prime}$.

### v2

@@ 5.1 Topology of links, spheres, and balls
Finally, we check that the actual simplicial complexes coincide. For the simplicial complex $\operatorname{lk}(\square,X)$, the set of vertices $\{\square^{\prime}_{0},\ldots,\square^{\prime}_{k}\}$ span a $k$-simplex if and only if the cubes $\{\square^{\prime}_{0},\ldots,\square^{\prime}_{k}\}$ are all contained in common larger cube $\square^{\prime\prime}$. Considering $\operatorname{lk}(v,X)$, this collection of cubes correspond to a collection of simplices $\{\tau^{\prime}_{0},\ldots,\tau^{\prime}_{k}\}$ in $\operatorname{lk}(v,X)$ all containing $\tau$ as a codimension one face. The larger cube $\square^{\prime\prime}$ leads to a larger simplex $\tau^{\prime\prime}\subset\operatorname{lk}(v,X)$ containing all of the $\tau^{\prime}_{i}$ as faces. Viewing the $\tau^{\prime}_{i}$ as a collection of vertices in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$, it follows that they span a simplex in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$. This shows that simplices in $\operatorname{lk}(\square,X)$ give rise to simplices in $\operatorname{lk}(\tau,\operatorname{lk}(v,X))$. One can reverse this argument to show the converse. Thus the bijection on vertex sets described above also bijectively identifies simplices of the two complexes, establishing that the two simplicial complexes are isomorphic. ∎
@@
Next we turn our attention to the topology of the spherical sets $S(\sigma,\Sigma)$. Recall that this set is a small $\epsilon$-sphere centered around the simplex $\sigma$, and it inherits a natural polyhedral structure from the ambient simplicial complex $\Sigma$. Indeed, each point $p\in S(\sigma,\Sigma)$ is contained in the interior of a unique simplex in $\Sigma$ that splits as a join $\sigma_{p}*\tau_{p}$ where $\sigma_{p}\subset\sigma$ is a face of $\sigma$ and $\tau_{p}$ is a simplex disjoint from $\sigma$. The intersection of $S(\sigma,\Sigma)$ with the simplex $\sigma_{p}*\tau_{p}$ is a copy of the polyhedron $\sigma_{p}\times\tau_{p}$, which we can think of as the $\epsilon$-slice” $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$ in the join:
@@
Now given $\tau$, express it as a join $v*\tau^{\prime}$ where $v\in\hat{S}(\sigma,\Sigma)$ is a vertex and $\tau^{\prime}\subset\hat{S}(\sigma,\Sigma)$ is a codimension one face. By the inductive hypothesis, there exists a vertex $w^{\prime}\subset\sigma$ with the property that $w^{\prime}*\tau^{\prime}$ is a simplex in $\Sigma$. If $v$ is connected to $w^{\prime}$, then $\tau*w^{\prime}$ is a simplex since $\Sigma$ is flag, and we are done. If $v$ is not connected to $w^{\prime}$, then $v$ is connected to some vertex $v^{\prime}\in\sigma$, with $v^{\prime}\neq w^{\prime}$. For each vertex $w\in\tau^{\prime}$, consider the ordered $4$-tuple $\{v,v^{\prime},w^{\prime},w\}$. This defines a square in $\Sigma$, and we know by construction that $v$ is not connected to $w^{\prime}$. Then $5$-largeness forces $w$ to be connected to $v^{\prime}$. Since this holds for arbitrary $w\in\tau^{\prime}$, we conclude that all the vertices of $\tau^{\prime}$ are connected to $v^{\prime}$. Again, the fact that $\Sigma$ is flag implies that $\tau*v^{\prime}$ is a simplex, thus completing the proof. ∎
@@
[…]
@@
We first consider the topology of $\hat{S}(\sigma,\Sigma)$. There is a natural map $\rho\colon S(\sigma,\Sigma)\rightarrow\hat{S}(\sigma,\Sigma)$, given by projecting away from the simplex $\sigma$. More explicitly, on each simplex
@@
incident to $\sigma$, the intersection with $S(\sigma,\Sigma)$ is the polyhedral slice $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$, while the intersection with $\hat{S}(\sigma,\Sigma)$ is the top slice $\tau_{p}\cong(\sigma_{p}\times\tau_{p}\times\{1\})/\sim$. The map $\rho$ is then just the projection to $\{1\}$ in the interval coordinate. In terms of the polyhedral structure on $\hat{S}(\sigma,\Sigma)$, $\rho$ collapses the $\sigma_{p}$ factor in each polyhedron $\sigma_{p}\times\tau_{p}$. Note that $\rho$ is surjective by 5.3. Thus the intersection of the preimage under $\rho$ of any point with any given polyhedron of $S(\sigma,\Sigma)$ is contractible or empty.
@@
To verify the corresponding statement for $\hat{B}(\sigma,\Sigma)$, we note that $\hat{B}(\sigma,\Sigma)$ is the union of simplices of the form $\sigma_{p}*\tau_{p}$, where $\tau_{p}$ is a face in $\hat{S}(\sigma,\Sigma)$ and $\sigma_{p}$ is a face of $\sigma$ by 5.3. We now form an auxiliary space $B(\sigma,\Sigma)$ and a continuous quotient map $\phi\colon B(\sigma,\Sigma)\rightarrow\hat{B}(\sigma,\Sigma)$. Recall that the space $S(\sigma,\Sigma)$ is the $\epsilon$-sphere centered at $\sigma$, identified with the $\epsilon$-slice $\sigma_{p}\times\tau_{p}\times\{\epsilon\}$ in the join description of each simplex incident to $\sigma$. We let $B(\sigma,\Sigma)$ denote the closed $\epsilon$-ball centered at $\sigma$. Thus $B(\sigma,\Sigma)$ gets identified with $(\sigma_{p}\times\tau_{p}\times[0,\epsilon])/\sim$ in the join description of each simplex incident to $\sigma$; recall that at the 0-level we collapse the $\tau_{p}$ coordinates to points. Since we can contract the interval factors down, we clearly have that $B(\sigma,\Sigma)$ deformation retracts to $\sigma$, hence it is contractible.
The space $B(\sigma,\Sigma)$ contains a copy of $S(\sigma,\Sigma)$ at level $\epsilon$, and we now define $\phi$ to be the quotient map $\rho$ on the subset $S(\sigma,\Sigma)$. As we can stretch out the $0,\epsilon]$ coordinate to $[0,1]$, we see that the quotient can naturally be identified with $\hat{B}(\sigma,\Sigma)$. Moreover point preimages under $\phi$ are either points when $x\notin\hat{S}(\sigma,\Sigma)$ or coincide with the point preimage under $\rho$ when $x\in\hat{S}(\sigma,\Sigma)$, and hence point preimages are always contractible. Thus the map $\phi$ is cell-like, and hence it is a homotopy equivalence. We conclude that the space $\hat{B}(\sigma,\Sigma)$ is contractible, as desired. ∎
@@ 5.2 Analysis of combinatorial spheres in thickenings
We now return to analyzing the thickenings $T_{n}$ from Section 4 in order to prove 4.1. Adopting all notation from that section, recall that the goal is to prove that $\operatorname{cd}(S(\sigma,T_{n}))\leq n-1$ for all nonempty simplices $\sigma$ of $T_{n}$. Equivalently, we want to show that $\overline{H}{}^{i}(S(\sigma,T_{n}))=0$ for all $i\geq n$ and $\emptyset\neq\sigma\subset T_{n}$.
@@
To complete the proof of the lemma, it remains to verify property (3). Let $Z$ denote the union of all the cubes containing $\square_{\sigma}$ and $\hat{Z}$ be the cube subcomplex of $Z$ consisting of all cubes in $Z$ that do not intersect $\square_{\sigma}$. There is a natural homeomorphism $\hat{Z}\cong\square_{\sigma}\times\operatorname{lk}(\square_{\sigma},X)$, so $\hat{Z}$ is homotopy equivalent to $\operatorname{lk}(\square_{\sigma},X)$. From the definition of $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$, any simplex $\tau\subset\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ has the property that $\alpha(V(\tau))$ is disjoint from $\square_{\sigma}$, and hence it lies in V(Z^)V($\hat${Z}). In other words, the entire complex $\operatorname{lk}^{\prime\prime}(\sigma,\operatorname{Th}_{\alpha}(X))$ lies “above” the cube complex $\hat{Z}$.
[…]
@@
[…]
@@
The bottom of the induction is the case of a cube $\square\subset Z$ with $\dim(\square)=\dim(\square_{\sigma})+1$. In this case, $\square=\square_{\sigma}\times I$, and $\hat{Z}\cap\square$ is the face of $\square$ opposite $\square_{\sigma}$. This forces the two subcomplexes above $\hat{Z}\cap\square$ to coincide, and hence we can take the identity map as the deformation retraction. Now for the inductive step, we assume that we are already provided the deformation retractions for the portions of the simplex that lie above the relevant lower dimensional faces of the cube $\hat{Z}\cap\square$. We would like to extend the deformation retraction to the rest of the simplicial complex lying above $\hat{Z}\cap\square$.
@@
[…]
@@
$\operatorname{cd}(\operatorname{lk}(\sigma,T_{n}))=\operatorname{cd}(\operatorname{lk}(\tau,T_{n-1}))\leq n-2$
@@
Note that every vertex $v$ in $\hat{S}(\sigma,T_{n})$ is at distance one from some vertex of $\sigma$ by definition. For such a vertex $v$, we can look at all the vertices of $\sigma$ that are adjacent to $v$. This defines a face of $\sigma$, and gives us a map $\pi\colon V(\hat{S}(\sigma,T_{n}))\rightarrow\mathcal{F}(\sigma)$ from the set of vertices in $\hat{S}(\sigma,T_{n})$ to the set $\mathcal{F}(\sigma)$ of faces of $\sigma$. We think of $\pi(v)$ as encoding the face of $\sigma$ that is closest to the vertex $v$. We can now partition the vertices of $\hat{S}(\sigma,T_{n})$ according to their image under $\pi$; see Figure 8. Next, use this partition to form a filtration of $\hat{S}(\sigma,T_{n})$ by subcomplexes. Denote by $S_{i}$ the subcomplex of $\hat{S}(\sigma,T_{n})$ spanned by the preimage under $\pi$ of all faces of codimension $\leq i$. We clearly have containments $S_{0}\subset S_{1}\subset\cdots\subset S_{k}=\hat{S}(\sigma,T_{n})$, where $k=\dim(\sigma)$. This gives us a filtration of the complex $\hat{S}(\sigma,T_{n})$ by the subcomplexes $S_{i}$; again, see Figure 8.
[…]
Figure 8: The top row shows $\pi\colon V(\hat{S}(\sigma,T_{n}))\rightarrow\mathcal{F}(\sigma)$ for faces of $\sigma$ in each codimension. The bottom row shows the filtration $S_{0}\subset S_{1}\subset\cdots\subset\hat{S}(\sigma,T_{n})$.
@@
By construction, the vertices in $L_{\tau}$ can be identified with vertices in $\operatorname{lk}(\tau,T_{n})$ by viewing the link as a subcomplex in the flag complex $T_{n}$. We claim that the vertices of $\operatorname{lk}(\tau,T_{n})\setminus\{L_{\tau}\cup\hat{\sigma}\}$} are precisely the vertices of $\operatorname{lk}(\tau,T_{n})$ at combinatorial distance one from $\hat{\sigma}$. Indeed, if $v$ is at combinatorial distance one from $\hat{\sigma}$, this means there is some vertex $w\in\hat{\sigma}$ connected by an edge to $v$. In terms of the embedding of $\operatorname{lk}(\tau,T_{n})$ into $T_{n}$, the simplex $\hat{\sigma}\subset\operatorname{lk}(\tau,T_{n})$ corresponds to the face of $\sigma$ opposite $\tau$, so $w\in\sigma\setminus\tau$. Since $v\in\operatorname{lk}(\tau,T_{n})$ is connected by an edge to $w$, we have that $\tau\cup\{w\}\subseteq\pi(v)$. This forces the codimension of $\pi(v)$ to be at most i1i-1, which means that $v\in S_{i-1}$. This argument can be reversed to show that the vertices in $L_{\tau}$ are precisely the vertices in $\operatorname{lk}(\tau,T_{n})$ whose combinatorial distance to $\hat{\sigma}$ is at least two. Of course, the vertices of $\hat{\sigma}$ in $\operatorname{lk}(\tau,T_{n})$ are precisely those in the face of $\sigma$ opposite $\tau$, so they do not lie in $S_{i-1}$. We conclude that the set $\partial L_{\tau}$ consists of precisely the vertices in $\operatorname{lk}(\tau,T_{n})$ that are exactly at distance $1$ from $\hat{\sigma}$.
@@
Consider a pair of simplicial complexes $Z,Z^{\prime}$ that can be expressed as gluings $Z=A\cup_{C}B$ and $Z^{\prime}=A\cup_{C}B^{\prime}$, where each inclusion of $C$ into $A$ induces the same maps on reduced cohomology. Assume $B$ is contractible, $\operatorname{cd}(Z)\leq n-1$, and $\operatorname{cd}(B^{\prime})\leq n-1$. Then $\operatorname{cd}(Z^{\prime})\leq n-1$.
@@
[…]
@@
[…]
By hypothesis, we have that $\operatorname{cd}(B^{\prime})\leq n-1$, so $\overline{H}{}^{i}(B^{\prime})=0$ for $i\geq n$. In particular, when $i\geq n$ the sequence above reduces to:
[…]
@@
We want to show that $\operatorname{cd}(S(\sigma,T_{n}))\leq n-1$ for every nonempty simplex $\sigma\subset T_{n}$. From 5.4, we have a homotopy equivalence $S(\sigma,T_{n})\simeq\hat{S}(\sigma,T_{n})$, so we instead show that $\operatorname{cd}(\hat{S}(\sigma,T_{n}))\leq n-1$. As discussed above, the set $\hat{S}(\sigma,T_{n})$ has a filtration
@@
We will consider the cohomological dimension of each stage in this filtration, and check that it always remains less than or equal to n1n-1.
@@
  • $B$ is the subcomplex of $Z$ consisting of all simplices incident to $\hat{\sigma}$ (from 5.10), which coincides with $\hat{B}(\hat{\sigma},Z)$, and is thus contractible by 5.4, since $T_{n}$, hence $Z$, is flag and 55-large;
@@
  • $C\coloneqq\partial L_{\tau}\subseteq S_{i-1}$ from the conclusions about $\partial L_{\tau}$ in 5.10, hence $C$ is contained in $B$ and $B^{\prime}$.
@@
Declarations of Competing Interests
The authors have no competing interests to declare that are relevant to the content of this article.
Data Availability Statement
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.